OpenStudy (anonymous):

how do you prove 2sin^2(t)=2-3sin(2t)-4cos^2(t)

OpenStudy (anonymous):

so far i have 2-2cos^2t = 2-6sintcost-4cos^2t

OpenStudy (anonymous):

({1-\cos(2t)} =

OpenStudy (anonymous):

\[({1-\cos(2t))} =\] ... Need to fix your identity on the left.